In a box , there are 3 red marbles,4 green marbles and 5 yellow marbles.in how many ways can you draw three marbles if 1 they are of the same color
2)they are of 3 different colors?

Answer by math_helper(2461)

I'm pretty sure this was answered already, but here it is again.
1. Order doesn't matter so we use the combination formula C(n,k) = n!/((n-k)!k!). The outcomes in this case are also mutually exclusive, so we simply add them:
C(3,3) + C(4,3) + C(5,3) = 1 + 4 + (5*4/2) = 15

2. Order doesn't matter, so another combination scenario. Use counting principle (independent events multiply the number of possible outcomes):
C(3,1)*C(4,1)*C(5,1) = 3*4*5 = 60